# Diff-PIC: Revolutionizing Particle-In-Cell Simulation for Advancing Nuclear Fusion with Diffusion Models

### v2

@@ Abstract
Sustainable energy is a crucial global challenge, and recent breakthroughs in nuclear fusion ignition underscore the potential of harnessing energy extracted from nuclear fusion in everyday life, thereby drawing significant attention to fusion ignition research, especially Laser-Plasma Interaction (LPI). Unfortunately, the complexity of LPI at ignition scale renders theory-based analysis nearly impossible – instead, it has to rely heavily on Particle-in-Cell (PIC) simulations, which is extremely computationally intensive, making it a major bottleneck in advancing fusion ignition. In response, this work introduces Diff-PIC, a novel paradigm that leverages conditional diffusion models as a computationally efficient alternative to PIC simulations for generating high-fidelity scientific data. Specifically, we design a distillation paradigm to distill the physical patterns captured by PIC simulations into diffusion models, demonstrating both theoretical and practical feasibility. Moreover, to ensure practical effectiveness, we provide solutions for two critical challenges: ❶ We develop a physically-informed conditional diffusion model that can learn and generate meaningful embeddings for mathematically continuous physical conditions. This model offers algorithmic generalization and adaptable transferability, effectively capturing the complex relationships between physical conditions and simulation outcomes; and ❷ We employ the rectified flow technique to make our model a one-step conditional diffusion model, enhancing its efficiency further while maintaining high fidelity and physical validity. Diff-PIC establishes a new paradigm for using diffusion models to overcome the computational barriers in nuclear fusion research, setting a benchmark for future innovations and advancements in this field.
@@ 1 Introduction
Sustainable energy stands as one of the paramount challenges of our era. The recent successful demonstration of fusion ignition [2] underscores the transformative potential of fusion as a sustainable energy source. In 2023 and 2024, the National Ignition Facility (NIF) achieved groundbreaking milestones, generating 3.4 MJ and 5.2 MJ of fusion energy from input energies of 2.2 MJ, respectively. These substantial energy gains emphasize the necessity for a deeper understanding of the fundamental sciences to enhance ignition efficiency. However, the complex and nonlinear nature of Laser-Plasma Interaction (LPI) in physics poses significant challenges for theory-based analysis, necessitating a heavy reliance on Particle-in-Cell (PIC) simulations [64, 35, 61, 5, 39]. Despite being the preeminent standard for modeling the physics of LPI, PIC simulations are exceedingly computationally intensive, often requiring tens of millions of CPU hours and consuming millions of dollars for high-resolution outputs [22, 16, 7]. The computational overhead of PIC simulations has become a bottleneck in fusion research, hindering progress toward the development of practical fusion power plants. Moreover, it is paradoxical that these extensive simulations, which aim to address environmental energy challenges, exacerbate the carbon footprint. Consequently, there is an urgent need for the development of innovative methodologies capable of producing high-quality data with substantially reduced computational burden.
@@
Over the years, numerous CPU-GPU optimizations have been developed to maximize computational efficiency. While these efforts are invaluable, they do not sufficiently address the inherent computational overhead. Recent advancements in generative AI, particularly diffusion models, present a novel approach to these challenges. Diffusion models [56, 24, 59] have demonstrated exceptional capabilities in Computer Vision (CV) [58, 49, 25, 50, 43, 52], synthesizing highly complex and high-dimensional data distributions that match real data with high fidelity. This has sparked significant interest in their potential for generating scientific synthetic data, as recent applications of diffusion models in molecular dynamics simulations have demonstrated their promise in this domain [69, 46].
Theoretically, by distilling the intricate physical patterns captured by PIC simulations into score functions for diffusion models, these models could potentially synthesize complex and high-resolution PIC simulations. In practice, however, the effectiveness of this approach to alleviate the current bottleneck in fusion research remains uncertain. Achieving this goal demands, first of all, that the synthesized data maintain physical soundness, and secondly, that there is an acceleration in simulation speed compared to conventional PIC methods. Consequently, two critical challenges naturally arise: ❶ PIC simulations utilize physical conditions as inputs, which differ significantly from the typical inputs in CV tasks, such as text prompts or specific classes. These physical conditions require continuity, meaning the synthesized results must vary smoothly with changes in input conditions. ❷ Diffusion models generally demand multiple denoising steps to produce high-fidelity data. In this context, it is essential to consider the speedup that diffusion models can provide over traditional PIC simulations while preserving high fidelity.
@@
In light of this perspective, this paper aims to offer the community an alternative to PIC simulations for generating synthetic LPI data and to tackle the two challenges mentioned above. Specifically, we propose an effective and efficient distillation paradigm, titled Diff-PIC (see Fig. 1), that leverages the entire denoising process of diffusion models to generate any arbitrary snapshot produced by PIC simulations. This paradigm treats each snapshot of the system’s evolution as a unique data distribution, thereby enhancing systemic efficiency. To address the first challenge ❶, we develop a conditional diffusion model with a physically-informed condition encoder. This encoder allows the conditional diffusion model to learn the relationship between continuous input conditions and generated data, making the physical phenomenon captured by PIC simulation easily distilled to Diff-PIC. For the second challenge ❷, we employ the rectified flow technique to further optimize the runtime efficiency of the proposed conditional diffusion model, demonstrating orders-of-magnitude speedup compared to PIC simulations.
In summary, our work represents the first known effort to tackle the imperative challenges associated with generating high-quality PIC simulation data for LPI using diffusion models. The core contributions of our work include:
  • We propose Diff-PIC, a pioneering study that utilizes diffusion models as a computationally efficient alternative to PIC simulations for the generation of synthetic data in fusion ignition research (see §3.1). By making all resources publicly available, we aim to establish Diff-PIC as a robust baseline for future improvements and a valuable benchmark for the research community, thereby accelerating the advancement in synthetic data generation for fusion ignition research.
  • We develop a physically-informed conditional diffusion model (see §3.2) that seamlessly integrates physical simulation conditions into the diffusion model. The designed condition encoder facilitates the generalization of effective simulation data within and beyond existing simulation parameters, endowing the model with robust generalization capabilities and adaptable transferability.
@@
  • Experimental results in §4.2 demonstrate that our method achieves a remarkable speedup of 16,200 times compared to traditional PIC simulations while preserving high fidelity and physical validity of the generated data, e.g., 0.341, 34.3, 8.98e-5 in terms of FID, SWD, and MMD, respectively.
@@ 2 Background
Inertial Confinement Fusion (ICF) is a method of achieving nuclear fusion by using intense energy pulses to compress and heat small fuel pellets [28, 9], typically containing isotopes of hydrogen such as deuterium and tritium. This process unfolds the nuclear fusion reaction as delineated below:
${}^{2}\mathrm{H}+{}^{3}\mathrm{H}\rightarrow{}^{4}\mathrm{He}+{}^{1}\mathrm{n}+\textbf{Energy}.$ (1)
Given the ubiquity of these hydrogen isotopes in the ocean, there is immense potential to harness “near-infinite” energy through the study of nuclear fusion. In particular, the goal of the study is to achieve the conditions of temperature and pressure that are sufficient to initiate fusion reactions, and attain a positive net energy gain (i.e., output energy surpassing the input). To optimize the sophisticated initiation of fusion, an advanced understanding of processes is imperative, especially for LPI. For such purposes, PIC is considered a crucial tool to provide theoretical insights into LPI, due to its capability of predicting and interpreting physical phenomena.
@@
In essence, PIC is an iterative finite element method applied to atomic particles such as electrons and ions. Within each iteration, particles are systematically arranged into discrete cells according to their spatial distribution, with their positions and velocities being updated over infinitesimally small time steps, typically on the scale of femtoseconds (i.e., $10^{-15}$ seconds). In contrast to molecular dynamics widely applied in biotechnology [42, 21, 8, 3, 14], PIC simulations in LPI are characterized by the presence of intensive electromagnetic fields, which exert a significant influence on particle trajectories as follows:
@@
Consequently, simulating laser-plasma interactions at the scale of even hundreds of Picoseconds (i.e., $10^{-10}$ seconds) requires the execution of hundreds of thousands of sophisticated PIC iterations. This imposes significant demands on computational storage and processing capabilities. Moreover, the computational complexity of the PIC method can be as high as $N^{2}$ for $N$ particles, thereby impeding scalability and computational efficiency. As a result, the PIC methodology has emerged as a stringent bottleneck in fusion research, significantly constraining progress in this domain.
Diffusion models have emerged as a prominent class of generative models within the realm of artificial intelligence, offering an innovative methodology for the synthesis of high-fidelity data [24, 25, 49]. Diverging from traditional generative models such as Generative Adversarial Networks [23] and Variational Autoencoders [31], diffusion models are named after the concept of diffusion in physics. These models are grounded in the principles of stochastic processes, particularly exploiting the notion of diffusion to conceptualize data generation and transformation.
Diffusion models function on the principle of simulating a diffusion process that progressively transforms simple, structured data into a complex distribution and vice versa. The core idea involves two main processes: (1) The Forward Process (Diffusion). In this process, Gaussian noise is gradually added to the data over a series of time steps, transforming the original data distribution into a tractable Gaussian distribution. Mathematically, this can be represented as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\alpha_{t})\mathbf{I}),$ (3)
where $\alpha_{t}$ regulates the noise level at each time step $t$. (2) The Reverse Process (Denoising). The objective of the reverse process is to reconstruct the original data from the noisy distribution by iteratively denoising the data. This denoising procedure is typically parameterized by a neural network, which is trained to approximate the inverse of the forward diffusion process:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)),$ (4)
where $\mu_{\theta}$ represent the mean predicted by the neural network and $\Sigma_{\theta}$ are untrained time dependent constants. Given their capacity to learn and represent data distributions, diffusion models hold significant promise for synthesizing data derived from Particle-In-Cell (PIC) simulations with enhanced speed and energy efficiency.
@@ 3 Diff-PIC
This section introduces Diff-PIC, a novel physically-informed conditional diffusion model tailored for generating high-fidelity synthetic data for fusion ignition research. As illustrated in Fig. 2, this paradigm conceptualizes each arbitrary snapshot in PIC simulations as a sample from a separate distribution, utilizing the denoising mechanism of diffusion models to predict an arbitrary snapshot $\mathcal{X}(t_{as})$ within the system’s evolution. Subsequently, we delve into the Physically-Informed Condition Encoder, crafted to tackle the physic condition encoding challenge in this context. Lastly, the Rectified Flow Acceleration technique is employed, significantly accelerating the denoising process and thereby enhancing the efficacy of the proposed conditional diffusion model.
Figure 2: Pipeline of the proposed Diff-PIC.
3.1 The Overall Distillation Paradigm
By extending the power of diffusion models to PIC simulations, we aim to distill the physical phenomena captured by PIC simulations into diffusion models. This distilling process enables the diffusion model to learn and replicate the complex behaviors and patterns observed in PIC simulations, providing a powerful tool for generating high-fidelity synthetic data in nuclear fusion research.
Concretely, we propose a distillation paradigm as follows: under a specified condition set $\theta$, using the entire denoising process of diffusion models to predict an arbitrary snapshot $\mathcal{X}(t_{as})$ of the system’s evolution, each representing a unique data distribution. $t_{as}$ specifies the simulation time step, different from the denoising step $t$. Therefore, our goal is to learn:
$\mathcal{X}(t_{as})=U(\epsilon~{}|~{}t_{as},\theta).$ (5)
With $U$, a diffusion model utilizing the vanilla U-Net for mapping Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$ to $\mathcal{X}(t_{as})$ under the simulation conditions of $\theta$. Particularly, conditions $\theta$ encompass variables including “Electron Temperature,” “Ion Temperature,” and “Laser Intensity,” allowing the model to generate $\mathcal{X}(t_{as})$, detailing the dynamics of each particle, alongside their electric fields at $t_{as}$.
To train the diffusion model, we construct a dataset of PIC simulation snapshots spanning various simulation conditions. Once trained, the diffusion model can generate synthetic snapshots of the system’s evolution at any desired time and under different physical conditions. These generated snapshots can be used for multiple purposes, including data augmentation, parameter exploration, and studying the effects of different physical conditions. Furthermore, the diffusion model aids in the interpretation and visualization of complex plasma phenomena, providing valuable insights for researchers in the field of nuclear fusion. Essentially, the proposed techniques offers two advantages:
@@ 3.1 The Overall Distillation Paradigm
  • Data Dependency Relaxation. By treating different snapshots under various condition combinations as separate distributions, we can sparsely sample across a specified range of conditions (see §1(c)). This approach reduces data dependency for model training, enabling the model to efficiently learn from a limited number of simulations while generalizing across a wide array of scenarios.
@@
  • Systemic efficiency. Unlike PIC simulations, which generate data sequentially over time, the proposed diffusion model can directly produce data for any target snapshot (see Fig. 3). This non-sequential behavior allows for more efficient data generation and analysis (see Table 4), enabling researchers to focus on specific times of interest without needing to simulate the entire LPI process.
3.2 Physically-Informed Condition Encoder

### v3

# Diff-PIC: Revolutionizing Particle-In-Cell Nuclear Fusion Simulation with Diffusion Models

@@ Abstract
The rapid development of AI highlights the pressing need for sustainable energy, a critical global challenge for decades. Nuclear fusion, generally seen as an ultimate solution, has been the focus of intensive research for nearly a century, with investments reaching hundreds of billions of dollars. Recent advancements in Inertial Confinement Fusion have drawn significant attention to fusion research, in which Laser-Plasma Interaction (LPI) is critical for ensuring fusion stability and efficiency. However, the complexity of LPI upon fusion ignition makes analytical approaches impractical, leaving researchers depending on extremely computation-demanding Particle-in-Cell (PIC) simulations to generate data, presenting a significant bottleneck to advancing fusion research. In response, this work introduces Diff-PIC, a novel framework that leverages conditional diffusion models as a computationally efficient alternative to PIC simulations for generating high-fidelity scientific LPI data. In this work, physical patterns captured by PIC simulations are distilled into diffusion models associated with two tailored enhancements: (1) To effectively capture the complex relationships between physical parameters and corresponding outcomes, the parameters are encoded in a physically-informed manner. (2) To further enhance efficiency while maintaining high fidelity and physical validity, the rectified flow technique is employed to transform our model into a one-step conditional diffusion model. Experimental results show that Diff-PIC achieves 16,200$\times$ speedup compared to traditional PIC on a 100 picosecond simulation, with an average reduction in MAE / RMSE / FID of 59.21% / 57.15% / 39.46% with respect to two other SOTA data generation approaches.
@@ 1 Introduction
Sustainable energy stands as one of the paramount challenges of our era, particularly with the rapid advancement of AI. The recent successful demonstration of fusion ignition [2] underscores the transformative potential of fusion as a sustainable energy source. In 2023 and 2024, the National Ignition Facility (NIF) achieved groundbreaking milestones, generating 3.4 MJ and 5.2 MJ of fusion energy, respectively, from 2.2 MJ input energy. Given that the estimated output could reach $\sim$120 MJ [53], there is a growing demand for a deeper understanding of the fundamental science behind ignition efficiency, especially the physical mechanisms governing the interaction between the laser and the plasma emitted when the laser bombards the fuel pellet. However, the Laser-Plasma Interaction (LPI) is a complex multi-body problem, which traditionally relies on time-stepping method, particularly, Particle-in-Cell (PIC) simulations [56, 32, 52, 4, 37]. Despite being the preeminent standard for modeling the physics of LPI, PIC simulations are exceedingly intensive in computation, often requiring tens of millions of CPU hours, consuming millions of dollars in order to obtain meaningful outputs [20, 14, 6]. The computational overhead of PIC simulations has become a daunting bottleneck in fusion research, raising the pressing need for innovative methodologies capable of generating high-quality scientific data with substantially reduced computational burden.
@@
Over the years, numerous CPU-GPU implementations have been developed for PIC simulations [17, 9, 44]. While invaluable, these efforts remain within the scope of the time-stepping approach that iteratively executes over infinitesimal time intervals, falling short in addressing the inherent computational overhead of long-term simulations. Recent advancements in generative AI – diffusion models, however, present a novel approach to bypass the constraint. Diffusion models [47, 22, 50] have demonstrated exceptional capabilities in Computer Vision (CV) [49, 42, 23], synthesizing highly complex data distributions that match real data with high fidelity. From the perspective of energy-based models [34, 21], a diffusion model effectively constructs a highly complex energy field that governs the evolution of variables, analogous to the motion of particles in cells. In fact, diffusion models are rooted in the diffusion concept in physics, where particles move according to the energy field. This has sparked significant interest in their potential for generating scientific data, as recent applications of diffusion models in molecular dynamics simulations have demonstrated their promise in this domain [60, 41].
Although diffusion models exhibit outstanding compatibility for generating PIC simulation data, two critical research gaps must be addressed. ❶ Physical soundness must be ensured. In contrast to traditional PIC simulations that directly take continuous physical parameters as constraints, it remains unclear how diffusion models can effectively capture and distill complex physical patterns. ❷ Substantial efficiency improvement must be achieved. The step-by-step denoising process in diffusion models is computationally demanding. Although the requirement for infinitesimal time intervals has been relaxed, the process still presents significant challenges similar to those faced by time-stepping methods, limiting the practicality of diffusion models as advanced alternatives for PIC simulations.
@@
In light of these challenges, we propose a distillation framework for physical patterns, titled Diff-PIC, that leverages diffusion models to efficiently generate a snapshot of arbitrary time, under arbitrary simulation parameters within certain ranges. To specifically address the two challenges above: ❶ We develop a conditional diffusion model with a Physically-Informed Parameter Encoder. This encoder allows the model to capture the relationship between continuous physical parameters and PIC simulation data, distilling physical phenomena into Diff-PIC. ❷ We employ the rectified flow technique to eliminate the requirement for the multi-step denoising, further optimizing the runtime efficiency of Diff-PIC. As highlighted in Fig. 1, orders-of-magnitude speedup is achieved compared to PIC simulations and other generative approaches including Generative Adversarial Networks (GAN) [3] and Normalizing Flow (NF) [62], dubbed "GAN-PIC" and "NF-PIC".
In summary, our work represents the first known effort to tackle the imperative challenges associated with generating high-quality PIC simulation data for LPI using diffusion models renowned in the field of generative AI. The core contributions of our work include:
  • We propose Diff-PIC, a pioneering study that utilizes diffusion models as a computationally efficient alternative to PIC simulations (see §3.1). By making all resources publicly available, we aim to establish Diff-PIC as a robust baseline and a valuable benchmark for the research community, thereby accelerating the advancement in scientific data generation for nuclear fusion research.
  • We develop a physically-informed conditional diffusion model (see §3.2) that seamlessly integrates physical simulation parameters into the diffusion model. The designed condition encoder facilitates the generalization of effective simulation data within and beyond existing simulation parameters, endowing the model with robust generalization capabilities and adaptable transferability.
@@
  • Experimental results in §4.2 demonstrate that our method achieves a remarkable speedup of 16,200 times compared to traditional PIC simulations while preserving high fidelity and physical validity of the generated data – 0.014, 0.019, 1.03 in terms of MAE, RMSE, and FID, respectively, with an average reduction of 59.21% / 57.15% / 39.46% compared to two other SOTA generative models.
@@ 2 Background
Inertial Confinement Fusion (ICF) is a method of achieving controlled nuclear fusion by using intense energy pulses to compress and heat small fuel pellets [27, 7], typically containing isotopes of hydrogen such as deuterium and tritium. This process unfolds the nuclear fusion reaction as delineated below:
${}^{2}\mathrm{H}+{}^{3}\mathrm{H}\rightarrow{}^{4}\mathrm{He}+{}^{1}\mathrm{neutron}+\textbf{Energy}.$ (1)
Given the ubiquity of these hydrogen isotopes in the ocean, nuclear fusion holds immense potential to provide "near-infinite" energy by achieving the necessary temperature and pressure conditions to initiate fusion reactions, attaining a positive net energy gain (i.e., output energy surpassing the input). To optimize the sophisticated initiation of fusion, an advanced understanding of the underlying LPI mechanism is essential. For this purpose, PIC is considered a crucial tool to provide theoretical insights into LPI, due to its capability of predicting and interpreting physical phenomena.
@@
In essence, PIC is an iterative time-stepping method applied to atomic particles such as electrons and ions. Within each iteration, particles are systematically arranged into discrete cells according to their spatial distribution, with their positions and velocities being updated over infinitesimally small time steps, typically on the scale of femtoseconds ($10^{-15}$ seconds). Unlike molecular dynamics widely applied in biology [19, 12], PIC simulations in LPI are characterized by intensive electromagnetic fields, which exert a significant influence on particle trajectories as follows:
@@
For a particle at position $\mathbf{r}_{i}$ with charge $q_{i}$ and mass $m_{i}$, the equation describes how the velocity $\mathbf{v}_{i}$ evolves according to an energy field that consists of electric field $\mathbf{E}$ and magnetic field $\mathbf{B}$. To ensure accuracy, simulating LPI at the scale of mere hundreds of picoseconds (i.e., $10^{-10}$ seconds) requires hundreds of thousands of sophisticated PIC iterations. This imposes significant demands on computational storage and processing capabilities. As a result, the PIC methodology has emerged as a stringent bottleneck in fusion research, significantly constraining progress in this domain.
Diffusion models have emerged as a prominent class of generative models within the realm of artificial intelligence, offering an innovative methodology for the synthesis of high-fidelity data [22, 23, 42]. Named after the physical concept, diffusion models use the idea of diffusion, which in physics refers to the random movement of particles from regions of high concentration to regions of lower concentration, often driven by thermal energy.
In machine learning, diffusion models are generative models that progressively add noise to data in a forward process, then gradually remove it in the reverse process to generate new data. From the perspective of energy-based models, the reverse process can be seen as moving variables through an energy landscape, where the model transitions variables from high-energy, noisy states (where data are unstructured) to low-energy, clean states (where data are structured). In this sense, diffusion models generate data by allowing variables to evolve in a manner akin to particles diffusing within an energy field, granting the opportunity to effectively apply physical methods to evaluate the diffusion, specifically, Langevin dynamics:
$d\mathbf{x}=-\nabla_{\mathbf{x}}E(\mathbf{x})dt+\sqrt{2\beta}d\mathbf{W_{t}}$ (3)
The evolution of variables $\mathbf{x}$ is influenced by both the energy field $E(\mathbf{x})$ and the stochastic motion term $\mathbf{W_{t}}$ regulated by the diffusion coefficient $\beta$. Through the dynamics, variables move towards low energy regions, conceptually mirroring the behavior of particles in PIC simulations. This similarity encourages an exploration of diffusion models as potential alternatives to traditional PIC simulations.
@@ 3 Diff-PIC
In this section, we introduce Diff-PIC, a physically-informed conditional diffusion model tailored for generating high-fidelity synthetic data for LPI in nuclear fusion. As illustrated in Fig. 2, accepting physical parameters as inputs, the parameters are encoded and integrated with the model main body through the physically informed parameter encoder. Additionally, the rectified flow mechanism is adopted to break free from the time-stepping paradigm in both PIC simulation and traditional diffusion models, unlocking the model’s capability to generate scientific data in one single step.
Figure 2: Workflow of the proposed Diff-PIC compared to traditional step-by-step PIC simulation.
3.1 The Overall Distillation Framework
Since LPI is governed by the behavior of electromagnetic field that dominates plasma dynamics, the tasks specifically assigned to Diff-PIC are to generate high-fidelity electric fields as representative cases under various physical parameters. Due to the polarization of the input laser, the resulting electric field only oscillates in a 2D plane, making 2D field descriptions sufficient to capture primary features of LPI. This approach is also a standard practice in PIC simulations.
In this work, we focus on the following critical physical parameters in LPI: “Electron Temperature ($T_{e}$) in $keV$,” “Ion Temperature ($T_{i}$) in $keV$,” and “Laser Intensity ($I$) in $W/m^{2}$.” Additionally, to enable direct and versatile generation of the electric field at a specific simulation time (a snapshot), we include a user-defined parameter $t_{as}$ that represents the simulation time for an arbitrary snapshot. Through a standard learning process in diffusion models [22, 48], the relationships between the four parameters and the resulting electric field are captured, as the model learns to progressively transform Gaussian noise into realistic electric fields. Once trained, the model takes the four parameters as inputs and produce electric field snapshots $\mathbf{E}(t_{as},\theta)$ corresponding to the specified parameters $\theta=\{T_{e},T_{i},I\}$. These generated snapshots can be used for multiple purposes, including data augmentation, parameter exploration, and studying the effects of different physical parameters. Furthermore, the diffusion model aids in the interpretation and visualization of complex plasma phenomena, providing valuable insights for researchers in the field of nuclear fusion. Essentially, the proposed techniques offer two advantages:
@@ 3.1 The Overall Distillation Framework
  • Data Dependency Relaxation. We treat snapshots from various parameter combinations as distinct distributions. This approach decouples data dependency in model training, enabling the model to efficiently learn from individual snapshots while generalizing across a wide array of scenarios.
@@
  • Systemic efficiency. Unlike traditional PIC simulations that generate data sequentially over time, the proposed diffusion model can directly produce data for any target snapshot (see Fig. 4). This non-sequential behavior allows for substantially more efficient data generation and analysis (see Table 3), enabling researchers to focus on specific times of interest without needing to simulate the entire LPI process from the beginning.
3.2 Physically-Informed Parameter Encoder